let f(x)=9-x, g(x)=xto the second power +x h(x)=x-2 and solve for g(f(3))
Answer by stanbon(75887)

let f(x)=9-x, g(x)=xto the second power +x h(x)=x-2 and solve for g(f(3))
-----------------
g(f(3) = g(9-3) = g(6) = 6^2 + 6 = 42
==========================================
Cheers,
Stan H.